# Nonequilibrium steady states in multi-bath quantum collision models

### v2

@@ III.1 Steady State Properties for Non-Markovian Collision Models
Considering first $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ for setting I (blue dashed line) we see that the system monotonically approaches the expected Gibbs state at temperature $\beta=2$. The dynamics appears effectively Markovian, despite the presence of the additional intra-environment interactions. Increasing the intra-environment interactions close to the maximal value is necessary before clear signatures of non-Markovianity emerge, as captured by the red solid curves in Fig. 3 where $\delta\!=0.95\tfrac{\pi}{2}\!$. However, these signatures depend on the value of $\Delta t$. Taking a smaller interaction time we find that, even in the presence of such extremely strong intra-environment interactions, the system’s dynamics still appears Markovian (gray dot-dashed line). Indeed, this can be confirmed by explicitly calculating the BLP measure, Eq. (15), shown in the inset. We see that for a given value of $\Delta t$, there is a critical value of $\delta$ below which the BLP measure is zero, and hence the dynamics appears Markovian. These dynamical differences notwithstanding, for all parameter choices in setting I the system is invariably driven to the expected thermal state. This can be readily confirmed by performing the Markovian embedding and calculating the steady state of system and memory Bassano. The nature of the system-bath interactions ultimately leads to homogenisation and, as will be relevant in later discussions, an asymptotic product state between system and bath units GiovannettiPRA2018.
@@
It follows then that for sufficiently weak intra-environment interactions, aligning with parameter ranges where suitable measures of non-Markovianity fail to reveal any strong memory effects, settings I and II appear essentially equivalent at the level of the steady states insofar as that they yield the same thermal steady state. Setting I, by construction, will drive the system to the Gibbs state at inverse temperature $\beta$ for any value of $\Delta t>0$ and $0\!\leq\!\delta\!<\!\tfrac{\pi}{2}$ GiovannettiPRA2018. By contrast, in the presence of strong intra-environment interactions setting II drives the system to a NESS which is in general at a different temperature. The deviation of the NESS in setting II from the canonical thermal state becomes more significant as the degree of non-Markovianity is increased.
@@ III.2 Steady State Correlations
We now consider the behavior of system-environment correlations. To do this we must employ a suitable measure and, for our purposes, we choose the negativity based on the postive partial transpose criterion. For a two qubit system, it is defined BiPartite_Neg,
@@
where $\lambda_{neg,j}^{A(BC)}$ are the negative eigenvalues of the partially transposed density matrix of the bipartition AA-BCBC. The genuine tripartite entanglement of the three-qubit state can then be calculated by taking a geometric average of the three possible bipartitions sabin2008classification,
@@
It is worth reminding that the above is in stark contrast to setting I’s steady state, which is driven to a factorised state between $S$ and $A_{n}$ and therefore clearly there is no entanglement present in the steady state GiovannettiPRA2018. This can be intuitively understood due to the fact that for the considered interactions, setting I leads to homogenization and therefore in the steady state the system and all bath auxiliary units are all in identical states and, as the intra-environment collisions correspond to a partial swap, they do not change the states. As demonstrated by Fig. 5, this is no longer the case in setting II, where the system evidently cannot homogenize and therefore each collisions is, generally, a strongly entangling operation.
@@ IV Stochastic Heat Currents from Trajectories
Quantum collision models lend themselves naturally to the simulation of quantum trajectories when we consider performing measurements on the auxiliary units immediately after their interaction with the system Brun_CM_Trajectories. It is known that even for non-Markovian collision models, individual trajectories can be averaged to recover the unconditional dynamics Kretschmer_CompositeCM; Whalen_NMCM_Trajectories; cilluffo2021microscopic. In this section we will show how the unconditional heat current can be recovered from two-point measurements on the auxiliary qubits. The stochastic heat current after one complete collision is calculated by performing a projective measurement, with measurement operator $P_{j}$, in the energy eigen-basis of the auxiliary unit. This measurement is performed twice, first before the auxilliary unit has undergone any interaction with any other system, and again after it has completed all of its interactions. This procedure is commonly referred to as a two-point measurement (TPM) scheme. The post-measurement states of the auxiliary unit after the first and second measurements are given by,
@@ V Conclusions
We have examined the steady state properties of a two-level quantum system undergoing an open dynamics simulated using quantum collision models. In the Markovian case, we considered a simple thermalisation process which can be effectively modelled via a standard GKSL master equation and considered two settings that drive the system to the same steady state.
@@
As a benchmark we considered the typical situation, referred to as setting I, where the collisional bath consists of identical auxiliary units all prepared in the same thermal state. For the system-environment interactions given by Eq. (6), setting I is known to lead to thermalisation, leaving the system in equilibrium with the bath and, furthermore, the system and environment share no correlations in the steady state. By contrast, setting II involves two collisional baths, whose auxiliary qubits are prepared in either their ground or excited states, and the bath temperature is encoded through the system-bath couplings. As a main result we have demonstrated that despite both settings recovering the GKSL master equation for the system in the $\Delta t\!\to\!0$ limit, setting II necessarily drives the system to a nonequilibrium steady state (NESS), supported by non-zero steady state heat currents. By exploiting the flexibility of collision models, we examined the impact of non-Markovian dynamics on the resulting steady state, by introducing additional intra-environment interactions. Although setting I and setting II exhibit similar amounts of non-Markovianity, as quantified via the information back flow, we showed that the non-Markovian case of setting II results in the system being driven away from the expected thermal state. Although setting I always reaches the canonical temperature of the bath, the non-equilibrium steady state of setting II can reach a different temperature. Through a careful analysis of the tripartite and bipartite entanglement, we have established that the non-equilibrium character of the steady state is rooted in the ability of setting II to create and maintain strong quantum correlations. We have also demonstrated how the average heat current can be recovered in setting II using a two-point measurement scheme, even in the presence of non-Markovianity.
@@
Our results provide insight into the dynamics and thermodynamics of thermalisation and equilibration and highlight that seemingly equivalent dynamics at the level of the system of interest can correspond to fundamentally different physical processes. The versatility of the collision model framework allows them to be employed in the modeling of practical quantum devices and offers a promising avenue for future work. One possible direction is to model non-Markovian fermionic environments relevant to semiconductor quantum dot devices coupled to electronic circuits goldhaber1998kondo; shim2018numerical; barthelemy2013quantum; iftikhar2018tunable; pouse2023quantum; petropoulos2024nanoscale, where we expect memory effects and strong system-bath correlations to play an important role. This builds on recent work showing that non-Markovian collision models are a powerful and systematic method for treating bosonic environments Lacroix2024MakingQCMExact. Another possibility would be to use the framework of non-Markovian collisional trajectories for measurement-based feedback control in non-Markovian environments, for example, to see the effects of specific intra-environment interactions and the resulting environmental memory on the ability to prepare a given target state HarwoodCMFeedback.
@@ V.1 Continuous-time limit of setting II
Following similar steps as Refs. CICCARELLO20221; cusumano2022quantum we can show that GKSL master equation, Eq. (4) is recovered in the time-continuous limit also for setting II. We assume an arbitrary form for the system-auxiliary interaction
@@
where $g_{\nu}$ are coefficients, $A_{\nu}$ and $B_{\nu}$ are sets of system and auxiliary operators, respectively, and $\Delta\rho_{n}=\rho_{n}-\rho_{n-1}$. By expanding the unitary operator governing a single collision up to second order in $\Delta t$, it can be readily shown that a finite difference equation for the state of the system can be derived CICCARELLO20221
@@ V.2 Heisenberg system-environment interaction
For collision models following setting I, this type of system-environment interaction has been extensively considered ScaraniPRL2002; ruariMauro; Bassano. (One can also consider an incoherent SWAP operation, see for example Ref. karpat2025transient.). For setting I, it should be evident that this choice of interaction has no impact on the steady state that the system is driven to; the dynamics will invariably homogenize the system with initial state of the auxiliary units in both the Markovian and non-Markovian regimes. For auxiliaries again initialised in thermal states, i.e. Eq. (5), it known that despite the dynamics driving the system to the same thermal state, for this type of interaction we no longer recover the dynamics governed by Eq. (4cusumano2022quantum. Nevertheless, as our focus is on the steady-state properties our interest lies in examining the how this choice of interaction reveals more substantial differences between settings I and II.
An immediate and remarkable difference emerges even in the Markovian case; i.e. $\delta\!=\!0$. The analytical form of the steady state in setting II is readily accessible, although too cumbersome to provide any direct insight. Nevertheless, it is diagonal in the system’s energy eigenbasis and, since we are considering a simple two-level system, it can therefore be assigned an effective temperature, $\beta_{e}$, via Eq. (17). While the explicit formula of $g_{e}$ or $\beta_{e}$ is too cumbersome, to illustrate the idea, in the case of very small $\Delta t$ we have
@@ Acknowledgements.
RMcE acknowledges support from Universities Ireland through the North/South Postgraduate Scholarship. RMcE and AKM acknowledge funding from Research Ireland through grant 21/RP2TF/10019. SC acknowledges support from the John Templeton Foundation Grant ID 62422.

### v3

@@ III.1 Steady State Properties for Non-Markovian Collision Models
Considering first $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ for setting I (blue dashed line) we see that the system monotonically approaches the expected Gibbs state at temperature $\beta=2$. The dynamics appears effectively Markovian, despite the presence of the additional intra-environment interactions. Increasing the intra-environment interactions close to the maximal value is necessary before clear signatures of non-Markovianity emerge, as captured by the red solid curves in Fig. 3 where $\delta\!=0.95\tfrac{\pi}{2}\!$. However, these signatures depend on the value of $\Delta t$. Taking a smaller interaction time we find that, even in the presence of such extremely strong intra-environment interactions, the system’s dynamics still appears Markovian (gray dot-dashed line). Indeed, this can be confirmed by explicitly calculating the BLP measure, Eq. (15), shown in the inset. We see that for a given value of $\Delta t$, there is a critical value of $\delta$ below which the BLP measure is zero, and hence the dynamics appears Markovian. These dynamical differences notwithstanding, for all parameter choices in setting I the system is invariably driven to the expected thermal state. This can be readily confirmed by performing the Markovian embedding and calculating the steady state of system and memory Campbell et al. (2018). The nature of the system-bath interactions ultimately leads to homogenisation and, as will be relevant in later discussions, an asymptotic product state between system and bath units Cusumano et al. (2018a).
@@
It follows then that for sufficiently weak intra-environment interactions, aligning with parameter ranges where suitable measures of non-Markovianity fail to reveal any strong memory effects, settings I and II appear essentially equivalent at the level of the steady states insofar as that they yield the same thermal steady state. Setting I, by construction, will drive the system to the Gibbs state at inverse temperature $\beta$ for any value of $\Delta t>0$ and $0\!\leq\!\delta\!<\!\tfrac{\pi}{2}$ Cusumano et al. (2018a). By contrast, in the presence of strong intra-environment interactions setting II drives the system to a NESS which is in general at a different temperature. The deviation of the NESS in setting II from the canonical thermal state becomes more significant as the degree of non-Markovianity is increased.
@@ III.2 Steady State Correlations
We now consider the behavior of system-environment correlations. To do this we must employ a suitable measure and, for our purposes, we choose the negativity based on the postive partial transpose criterion. For a two qubit system, it is defined Peres (1996),
@@
where $\lambda_{neg,j}^{A(BC)}$ are the negative eigenvalues of the partially transposed density matrix of the bipartition AA-BCBC. The genuine tripartite entanglement of the three-qubit state can then be calculated by taking a geometric average of the three possible bipartitions Sabín and García-Alcaine (2008),
@@
It is worth reminding that the above is in stark contrast to setting I’s steady state, which is driven to a factorised state between $S$ and $A_{n}$ and therefore clearly there is no entanglement present in the steady state Cusumano et al. (2018a). This can be intuitively understood due to the fact that for the considered interactions, setting I leads to homogenization and therefore in the steady state the system and all bath auxiliary units are all in identical states and, as the intra-environment collisions correspond to a partial swap, they do not change the states. As demonstrated by Fig. 5, this is no longer the case in setting II, where the system evidently cannot homogenize and therefore each collisions is, generally, a strongly entangling operation.
@@ IV Stochastic Heat Currents from Trajectories
Quantum collision models lend themselves naturally to the simulation of quantum trajectories when we consider performing measurements on the auxiliary units immediately after their interaction with the system Brun (2002). It is known that even for non-Markovian collision models, individual trajectories can be averaged to recover the unconditional dynamics Kretschmer et al. (2016b); Whalen (2019); Cilluffo et al. (2021). In this section we will show how the unconditional heat current can be recovered from two-point measurements on the auxiliary qubits. The stochastic heat current after one complete collision is calculated by performing a projective measurement, with measurement operator $P_{j}$, in the energy eigen-basis of the auxiliary unit. This measurement is performed twice, first before the auxilliary unit has undergone any interaction with any other system, and again after it has completed all of its interactions. This procedure is commonly referred to as a two-point measurement (TPM) scheme. The post-measurement states of the auxiliary unit after the first and second measurements are given by,
@@ V Conclusions
We have examined the steady state properties of a two-level quantum system undergoing an open dynamics simulated using quantum collision models, critically assessing the differences between models that employ a single thermal bath and multiple pure-state baths. As a benchmark, we considered the Markovian case of a simple thermalisation process which can be effectively modelled via a standard GKSL master equation and we leverage the flexibility of collision models to extend this to the non-Markovian setting. We have also shown that the two-bath setup allows for non-Markovian quantum trajectories to be defined. The two-bath setup therefore combines the flexibility of collision models to introduce nontrivial memory effects, with the ability to resolve interpretable stochastic quantities via TPM schemes. Our work builds on recent studies that have considered the steady state properties of Markovian collision models Guarnieri et al. (2020); Prositto et al. (2025a, b) and extends them to the multi-bath and non-Markovian regime
@@
We considered the typical situation, referred to as setting I, where the collisional bath consists of identical auxiliary units all prepared in the same thermal state. For the system-environment interactions given by Eq. (6), setting I is known to lead to thermalisation, leaving the system in equilibrium with the bath and, furthermore, the system and environment share no correlations in the steady state. By contrast, setting II involves two collisional baths, whose auxiliary qubits are prepared in either their ground or excited states, and the bath temperature is encoded through the system-bath couplings. As a main result we have demonstrated that despite both settings recovering the GKSL master equation for the system in the $\Delta t\!\to\!0$ limit, setting II necessarily drives the system to a nonequilibrium steady state (NESS), supported by non-zero steady state heat currents. By exploiting the flexibility of collision models, we examined the impact of non-Markovian dynamics on the resulting steady state, by introducing additional intra-environment interactions. Although setting I and setting II exhibit similar amounts of non-Markovianity, as quantified via the information back flow, we showed that the non-Markovian case of setting II results in the system being driven away from the expected thermal state. Although setting I always reaches the canonical temperature of the bath, the non-equilibrium steady state of setting II can reach a different temperature. Through a careful analysis of the tripartite and bipartite entanglement, we have established that the non-equilibrium character of the steady state is rooted in the ability of setting II to create and maintain strong quantum correlations. We have also demonstrated how the average heat current can be recovered in setting II using a two-point measurement scheme, even in the presence of non-Markovianity.
@@
Our results provide insight into the dynamics and thermodynamics of thermalisation and equilibration and highlight that seemingly equivalent dynamics at the level of the system of interest can correspond to fundamentally different physical processes. The versatility of the collision model framework allows them to be employed in the modeling of practical quantum devices and offers a promising avenue for future work. One possible direction is to model non-Markovian fermionic environments relevant to semiconductor quantum dot devices coupled to electronic circuits Goldhaber-Gordon et al. (1998); Shim et al. (2018); Barthelemy and Vandersypen (2013); Iftikhar et al. (2018); Pouse et al. (2023); Petropoulos et al. (2024), where we expect memory effects and strong system-bath correlations to play an important role. This builds on recent work showing that non-Markovian collision models are a powerful and systematic method for treating bosonic environments Lacroix et al. (2024). Another possibility would be to use the framework of non-Markovian collisional trajectories for measurement-based feedback control in non-Markovian environments, for example, to see the effects of specific intra-environment interactions and the resulting environmental memory on the ability to prepare a given target state Harwood et al. (2023).
@@ V.1 Continuous-time limit of setting II
Following similar steps as Refs. Ciccarello et al. (2022); Cusumano (2022) we can show that GKSL master equation, Eq. (4) is recovered in the time-continuous limit also for setting II. We assume an arbitrary form for the system-auxiliary interaction
@@
where $g_{\nu}$ are coefficients, $A_{\nu}$ and $B_{\nu}$ are sets of system and auxiliary operators, respectively, and $\Delta\rho_{n}=\rho_{n}-\rho_{n-1}$. By expanding the unitary operator governing a single collision up to second order in $\Delta t$, it can be readily shown that a finite difference equation for the state of the system can be derived Ciccarello et al. (2022)
@@ V.2 Heisenberg system-environment interaction
For collision models following setting I, this type of system-environment interaction has been extensively considered Scarani et al. (2002); McCloskey and Paternostro (2014); Campbell et al. (2018) (we remark that one can also consider an incoherent SWAP operation, see for example Ref. Karpat and Çakmak (2025)). For setting I, it should be evident that this choice of interaction has no impact on the steady state that the system is driven to; the dynamics will invariably homogenize the system with initial state of the auxiliary units in both the Markovian and non-Markovian regimes. For auxiliaries again initialised in thermal states, i.e. Eq. (5), it known that despite the dynamics driving the system to the same thermal state, for this type of interaction we no longer recover the dynamics governed by Eq. (4Cusumano (2022). Nevertheless, as our focus is on the steady-state properties our interest lies in examining the how this choice of interaction reveals more substantial differences between settings I and II.
An immediate and remarkable difference emerges even in the Markovian case; i.e. $\delta\!=\!0$. The analytical form of the steady state in setting II is readily accessible, although too cumbersome to provide any direct insight. Nevertheless, it is diagonal in the system’s energy eigenbasis and, since we are considering a simple two-level system, it can therefore be assigned an effective temperature, $\beta_{e}$, via Eq. (17). While the explicit formula of $g_{e}$ or $\beta_{e}$ is too cumbersome, to illustrate the idea, in the case of very small $\Delta t$ we take a Taylor expansion to first order and find
@@ Acknowledgements.
RMcE acknowledges support from Universities Ireland through the North/South Postgraduate Scholarship. RMcE and AKM acknowledge funding from Research Ireland through grant 21/RP2TF/10019. This work was supported by the John Templeton Foundation (Grant IDs 62422 and 63626).
The data associated to this work is available from McElvogue (2026).